Find the legs for a 20, 70, 90 triangle.

Answer by cleomenius(959)

I believe we need the measurement for one of the sides to complete this example.
Since this is a 90 degree triangle, we can use the trig functions:
sin = opp/hyp
cos = adj/hyp
tan = opp/adj.
Cleomenius.